Circular No. 6666

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970508 M. Morris, N. Mastrodemos, B. Zuckerman, C. McCarthy, E. Becklin,

P. Lowrance, and R. Chary, University of California at Los Angeles; and C. Barnbaum, Space Telescope Science Institute, report: “We imaged the optical transient associated with GRB 970508 (cf. IAUC 6649, 6654, 6655) with the Keck I telescope (+ NIRC camera + K_{s} filter). The K_{s} magnitudes of the variable on May 13.25 and 16.25 UT were 18.43 ± 0.05 and 19.21 ± 0.06, respectively, corresponding to fluxes of 28.3 ± 1.3 and 13.8 ± 1.7 μJy. The gradient of the decline in this interval is similar to that reported in the r band in the preceding 3-day interval (IAUC 6658, 6660). The May 13.25 K_{s} flux is slightly above the contemporaneous r-band flux, giving an r − K_{s} color of +2.3 ± 0.2. In these brief (∼ 10-min) images, the limit on the K_{s} flux from any associated galaxy within a 1′′.2–2′′.4 annulus around the transient is 5 μJy.”

M. Donahue, K. C. Sahu, M. Livio, and L. Petro, Space Telescope Science Institute; J. T. Stocke, University of Colorado; P. Groot, University of Amsterdam (UoA); and J. van Paradijs, UoA and University of Alabama, write: “A 15-min exposure in the I band taken on May 9.15 UT at the Kitt Peak 4-m telescope was used to get relative photometry of the optical counterpart of GRB 970508 (IAUC 6654, 6655, 6658), yielding I = 20.4 ± 0.1, assuming the magnitude of the ‘standard’ star of Schaefer et al. (IAUC 6658) to be I = 18.4 ± 0.1.”

V485 CENTAURI A. Olech, Warsaw University Observatory, writes: “CCD observations

of the dwarf nova V485 Cen during its recent outburst, collected with the 1.3-m Warsaw telescope at Las Campanas Observatory as a subproject of the OGLE-2 project, indicate that this star belongs to the SU UMa group of cataclysmic variables. A 3-hr I-band observation on May 17.1 UT shows three well-developed typical superhumps with an amplitude of 0.25 mag and period 57.7 min. Thus, V485 Cen has the shortest known period among SU UMa-type stars, with a period as much as 25 percent shorter than previously known objects.”

SUPERNOVA 1997bp IN NGC 4680 Unfiltered CCD magnitude estimate by S. Moretti and S. Tomaselli,

Forlí, Italy: May 14.92 UT, 15.6 (GSC comparison stars).

1997 May 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green